# The Atacama Cosmology Telescope: DR6 Gravitational Lensing and SDSS BOSS cross-correlation measurement and constraints on gravity with the EG statistic

### v1

@@ VI.2.5 Consistency with CIB-deprojection approach
In Fig. 8 we show the angular cross-power spectra for the CIB-deprojected ACT CMB lensing map and the baseline map restricted to the same footprint. The two measurements closely match, as they are highly correlated. We test for consistency between the two measurements using a PTE value test, accounting for the correlation. For the cosmological analysis range, we find PTE values of 0.40 and 0.07 for ACT CMASS and ACT LOWZ respectively. In the extended validation range, we find PTE values of 0.63 and 0.30 respectively. All PTE values indicate consistency between the two foreground mitigation approaches.
VI.2.6 Consistency between North and South observations
The angular power spectra are calculated using the full CMASS and LOWZ maps. We check the measurements for consistency between the North and South patches of the galaxy observations (see Fig. 2). By showing that the angular cross-power spectra calculated for each separately are consistent we confirm that it is justified to analyze the full map together.
In Fig. 9 we show the result for the angular cross-power spectra calculated for the North and South patches separately. We evaluate the consistency between the two measurements via a PTE value test. Since the two angular cross-power spectra do not overlap in the area and our largest scales considered () are significantly smaller than the separation between the two patches we assume them to be uncorrelated. In the cosmological analysis range, we find PTE values of 0.66 and 0.97 for ACT CMASS and ACT LOWZ respectively. For the extended range, the PTE values are 0.50 and 0.96 respectively. In all cases, the PTE values indicate full consistency between the North and South patches. We note that the values for ACT LOWZ are marginally above 0.95 which indicates that they agree closer than expected in 95% of tests. Since we perform a large number of PTE tests we assume that this is a chance occurrence and evaluate if there is a higher-than-expected number of high PTE values among our PTE tests when summarizing the distribution of all of them in Section VI.2.7.
@@ VI.2.6 Consistency between North and South observations
VI.2.7 Summary of PTE statistics
Throughout this work, we report a large number of PTE tests. To evaluate if they overall represent a distribution within statistical expectations we consider the distribution of all PTE tests reported in the analysis.
We report a total of 40 tests. For an unbiased set of tests, each tenth between 0 and 1 should have a number of PTE values sampled from a Poisson distribution with a mean of 4 with a statistical variation of 2. If our distribution deviates from this it could indicate that the PTE value estimates are biased due to, for example, under- or overestimation of the uncertainty, or, for the cosmological tests, due to a disagreement with the GR-based predictions.
Among the reported PTE tests there are none with PTE values below 1% or above 0.99%. The lowest PTE value across all PTE tests is 0.04 for a null test discussed in Section VI.2.4. Overall we have 3 PTE values below 0.10 which is well within statistical expectations. We show a histogram of the PTE tests in bins of 10% in Fig. 10. For each bin in the histogram, the number of PTE values is within statistical expectations.
Overall the distribution of PTE values is consistent with expectations and shows the robustness of the analysis.
VI.3 Measurement results for cross-correlation of ACT and SDSS BOSS
@@
We show the measurements using ACT together with CMASS and LOWZ respectively. The top panels show the measurements with the baseline simulation-based error estimates. The white background range indicates the scales used for the cosmological analysis. The full range plotted is the extended validation range for which we also performed null and consistency tests. For comparison, the theory curve for our fiducial cosmology is also shown with dots indicating the binned measurement using the NaMaster coupling matrix. Note that this theory curve is not tuned to the amplitude of the measurement: the galaxy bias used is based on the amplitude of the angular auto-power spectrum of the galaxy data. In this work, we focus on constraining the statistic for which we directly use the measurements as part of the estimator, and the predicted theory values for are estimated from independent constraints on . In future work, the cross-correlation measurements we make available as part of this work can be consistently fitted with angular power spectra theory curves by varying the galaxy bias and all cosmological parameters. Specifically, we can visually see that for ACT CMASS the theory curve tuned to the measured galaxy auto-correlation amplitude appears slightly high compared to the cross-correlation measurement. This implies that when jointly fitting the cross-correlation one will likely find a lower best fit value, which quantifies the amplitude, than for the CMASS galaxy data alone. To evaluate whether this difference is significant or within statistical expectations careful theoretical modeling of the angular power spectra is needed in future work. We here provide the tested and validated cross-correlation measurements between ACT and CMASS that could be used for such an investigation. The ACT DR6 lensing data has already been analyzed in this way for combinations with galaxy data from unWISE [63] and similar investigations have been done for Planck lensing data [e.g. 71]. For this work, we focus on interpreting the measurement at linear scales through the statistic to interpret implications about gravity.
In the middle panels, the error for each bin is compared. The simulation-based estimate of the errors is closely consistent with the analytic estimate. This consistency also extends to the off-diagonal elements of the covariance as discussed in Section VI.1.3. We also show the uncertainties based on a jackknife approach which have larger statistical uncertainty and are a systematic overestimate. The jackknife errors are consistent with the other estimates within statistical expectations.
@@
Based on the measured angular cross-power spectra we derive constraints on the statistic for the ACT DR6 CMB lensing data combined with CMASS and LOWZ in Section VII. In the following, we combine the measurement with results based on Planck CMB lensing data.
VI.4 Combined cross-correlation measurement with ACT and Planck
@@
In Fig. 13 we show the results for combining the full baseline ACT SDSS and Planck SDSS angular cross-power spectra. We show the correlation with the CMASS sample on the left and with the LOWZ sample on the right. For each, we show the results using ACT CMB lensing discussed in Section VI.3 and the results using Planck SDSS presented in Wenzl et al. [26]. We can visually see that the two measurements are closely consistent with each other, especially on large scales, which is expected since they are strongly correlated. Finally, we show the results for the minimum variance combination of the two measurements. In the bottom panels, we show the SNR for each measurement as well as the combination. We again additionally show the SNR estimates using the theory curves (dashed black lines) which represent a fair comparison between the measurements.
VII EG measurement
VII.1 Measurement approach for EG
@@
In Section VII.1.1 we discuss the creation of correctly correlated simulated maps and the estimate of the cross-covariance between and .
We use estimates of and for the full galaxy maps. To quantify the potential systematic caused by differences in galaxy bias between the full sample and the ACT DR6 overlap we constrain the galaxy bias difference and derive an estimate for our systematic error budget in Section VII.1.2.
When calculating the estimate using Eq. 20 we account for the full covariance of and , including their cross-covariance as shown in Fig. 3. To do this we calculate the ratio distribution of . Then in a second step, we calculate the ratio distribution of this estimate and resulting in a probability density function (PDF) for marginalized across the cosmological range or for individual bins . Notably, while we assume the measurements of and to be Gaussian distributed we do not assume the ratio to be Gaussian as the PDF is expected to be significantly asymmetric [26]. We report the median and the central confidence range of the PDF which generalize the best fit and uncertainty for the Gaussian case. For a full description of this calculation and a derivation of the PTE value tests to evaluate the results for consistency with scale independence () and for consistency with a predicted value () see Wenzl et al. [26].
VII.1.1 Cross-covariance estimation
@@
For our baseline result using ACT and SDSS BOSS, we leverage the set of 400 simulated map realizations to estimate the full covariance. As described in Section VI.1.2, we have 400 simulated map realizations with the CMB lensing map and two galaxy maps for each with one galaxy map for the angular auto-power spectrum and the other for the angular cross-power spectrum where an additional redshift dependent weight was applied. For each realization, the three maps are correlated as expected for our fiducial cosmology and they contain the same noise level as our measurements. Using Eq. 34 we estimate the simulation-based covariances of and individually as well as the cross-covariance between them. For the full analytic covariance, we use NaMaster as described in Section VI.1.2. For the covariance of we use a shot noise of and for we use as given by Eq. 30. For the cross-covariance, we use . The resulting full simulation-based and analytic covariances are visualized as a correlation matrix in Fig. 3. We find consistent results between the two methods within the statistical uncertainty of the simulation-based covariance. For the same bin, the correlation between the two measurements reaches around 33% for ACT + CMASS and 28% for ACT + LOWZ and drops off for smaller scales.
VII.1.2 Effective galaxy bias difference in ACT overlap
@@
We can perform a test for the ratio of the inferred between the overlap and the full sample. Assuming the underlying matter power spectra and redshift distribution are the same for both the full sample and the overlap with ACT, this ratio directly tests if the effective linear galaxy bias is consistent between the two. By assuming that a constant linear bias describes the data well in our cosmological range, and therefore , we can put a constraint on the fractional difference in bias between the two . Since this approach avoids having to marginalize over uncertainty in the matter power spectrum this gives us a strong constraint on the fractional bias difference of the bias between the samples.
@@
Based on this we conclude that while the ACT data does not cover the full SDSS footprint the effective galaxy bias of the SDSS samples is sufficiently uniform so that within our statistical constraining power we can combine the cross-correlation with the auto-correlation and RSD analysis for the SDSS datasets to measure . For completeness, we also will show the result when restricting the to the overlap with ACT (see Section VII.3).
VII.2 Baseline EG results
@@
In Fig. 15 we show the measurement of as a function of scale. These are calculated using the estimator of Eq. 20, using the measured angular cross-power spectra between the ACT DR6 lensing map and the SDSS BOSS DR12 CMASS and LOWZ galaxy samples together with the angular auto-power spectra of the galaxy samples and from an RSD analysis of the galaxy samples. We account for the cross-covariance between the cross- and auto-power spectra and calculate the PDF for without assuming Gaussianity in the result. The Figure shows the median of the PDF for each bin in angular scale and the uncertainty is based on the central 68.27% confidence interval of the PDF. The constraints are shown for three different estimates of the uncertainty. The leftmost in orange are the results with the simulation-based covariance, in the middle in blue the constraints using the analytic covariance, and on the right in green the constraints using a jackknife-based covariance. In Section VI.1.3 we described in detail how the simulation-based covariance and analytic covariance are consistent within statistical expectations. Here they give consistent results for within the statistical accuracy of the simulation-based covariance. The jackknife-based covariance serves as a consistency check. It is a systematic overestimate of the uncertainty of 20-30% and has large statistical uncertainty as it is based on only a small number of subsamples. Its advantage is that it is based on the data itself and could reveal additional sources of uncertainty not captured in the other methods. We here again find it to be consistent with the other methods within statistical expectations giving confidence to the accuracy of the approach. Finally Fig. 15 also shows the CDM prediction for reference. A key feature of the GR-based predictions for is that on linear scales the statistic is expected to be scale-independent. Therefore, before comparing the overall amplitude with the CDM-based prediction we can more generally test if the measurement is consistent with scale independence. We perform a PTE value test for consistency of the measurement with a scale-independent constant. The constant here is a free parameter not fixed to the CDM prediction. We find for ACT + CMASS and for ACT + LOWZ and therefore no statistical evidence for a deviation from scale-independence within the accuracy of the measurement. For ACT + CMASS 9 out of the 11 bins are below the GR prediction. However, the different bins are significantly correlated so to evaluate if this is statistically significant we combine the measurements under consideration of the full covariance.
@@
Since the measurement is consistent with a constant we combine the constraints across the cosmological analysis range into an overall constraint on . For this, we account for the full covariance between and shown in Fig. 3. For the jackknife-based uncertainty estimate, there are not enough samples to invert this full covariance, therefore these combined constraints are calculated using the simulation-based and the analytic covariances. Fig. 16 shows the PDFs for the overall constraints. The statistic is measured for the effective redshift of the galaxy sample as given by Eq. 25, which results in for CMB lensing combined with CMASS and in for combinations with LOWZ. At these redshifts the expected values for the statistic for our fiducial cosmology assuming CDM are given by and . These predictions are shown as black vertical lines in Fig. 16. The Gaussian PDF for the predictions is too highly peaked to show in the same plot so the uncertainty for the predicted values is shown as a gray band instead. These predictions can be compared with the measurements based on ACT + CMASS and ACT + LOWZ. In each case, the PDF using the simulation-based covariance and the analytic covariance are shown. The results using either covariance are consistent. For the simulation-based covariance, the median is highlighted with a vertical orange line, and the central 68.27% confidence interval is shaded in orange. These result in and respectively. We note that the PDFs, especially for the LOWZ case, are significantly non-Gaussian and result in asymmetric uncertainties. The asymmetry is such that the error towards higher values is larger than towards lower values.
@@
We test the measured values for consistency with the CDM prediction resulting in for ACT + CMASS and for ACT + LOWZ. Both measurements are within statistical expectations, showing no statistical deviation from the prediction based on CDM. For ACT + CMASS the 68.27% confidence range does not include the prediction, i.e. the measurement differs by more than the equivalent of from the prediction. However, the deviation is not statistically significant with , meaning the 95% confidence interval, approximately equivalent to the error of a Gaussian, includes the prediction. Therefore within the statistical constraining power of the measurements, the data shows no evidence for a deviation from the CDM prediction. The results and PTE value tests are summarized in Table 2. In Section VII.3, we evaluate if these conclusions are sensitive to analysis choices. For an overview of values in alternative gravity scenarios see Pullen et al. [28].
@@
In addition to the statistical uncertainty we also estimate the systematic bias in the results. We combine in quadrature the conservative systematic bias estimates from galaxy bias evolution with redshift (Section III.2), the level at which we validated the analysis pipeline (Section VI.2.1) and the upper limit on the effective galaxy bias difference between the full galaxy maps and the overlap with ACT DR6 lensing. We find systematic error budgets of 3% for ACT + CMASS and 4% for ACT + LOWZ. These are significantly smaller than the statistical uncertainties of the measurement (approximately 18% and 26% respectively) and therefore negligible when evaluating the consistency of the measurement with GR predictions.
VII.3 Sensitivity to analysis choices

### v2

@@ V.2.5 Consistency with CIB-deprojection approach
We calculated the angular cross-power spectra for the CIB-deprojected ACT CMB lensing map and the baseline map restricted to the same footprint. We found that the results closely match. We test for consistency between the two measurements using a PTE value test, accounting for the correlation. For the cosmological analysis range, we find PTE values of 0.40 and 0.07 for ACT CMASS and ACT LOWZ respectively. In the extended validation range, we find PTE values of 0.63 and 0.30 respectively. All PTE values indicate consistency between the two foreground mitigation approaches.
V.2.6 Consistency between North and South observations
SDSS has two separate patches on the sky, in the North and the South. The two patches are often observable in different seasons and include different latitudes. Higher latitudes can have less foreground contamination and the patches could have different calibrations, seasonal variations, and selection functions [72]. In this analysis we proceed analogous to the approach for the auto-correlation developed in [27], to test if, for the cross-correlations measured in this work, there are any statistically significant differences between the patches that might indicate that systematic effects are relevant. If no such differences are found it motivates an analysis of the data as a combined whole rather than two distinct patches.
The angular power spectra are calculated using the full CMASS and LOWZ maps. We check the measurements for consistency between the North and South patches of the galaxy observations (see Fig. 1). We find that both North and South patch correlations agree statistically.
In Fig. 6 we show the result for the angular cross-power spectra calculated for the North and South patches separately. We evaluate the consistency between the two measurements via a PTE value test. Since the two angular cross-power spectra do not overlap in the area and our largest scales considered () are significantly smaller than the separation between the two patches we assume them to be uncorrelated. In the cosmological analysis range, we find PTE values of 0.66 and 0.97 for ACT CMASS and ACT LOWZ respectively. For the extended range, the PTE values are 0.50 and 0.96 respectively. In all cases, the PTE values indicate full consistency between the North and South patches. We note that the values for ACT LOWZ are marginally above 0.95 which indicates that they agree closer than expected in 95% of tests. Since we perform a large number of PTE tests we assume that this is a chance occurrence and evaluate if there is a higher-than-expected number of high PTE values among our PTE tests when summarizing the distribution of all of them in Section V.2.7.
@@ V.2.6 Consistency between North and South observations
V.2.7 Summary of PTE statistics
Throughout this work, we report a large number of PTE tests. If a measurement is consistent with the null hypothesis then the PTE value by definition should be a sample from a uniform distribution between 0 and 1. When reporting many PTE tests the probability of detecting outliers increases. Under the assumption that the distinct reported PTE values are independent, we can analyze whether together they follow this uniform distribution. If they do it indicates that the statistical tests performed are unbiased and show consistency between the measurement and our null hypothesis. If they are not consistent it could point to issues with the test, the analysis pipeline, or an inconsistency of the measurement with GR. We here analyze both the overall distribution and check for stronger outliers that would be expected for the number of tests we perform.
We report a total of 40 tests. For an unbiased set of tests, each tenth between 0 and 1 should have a number of PTE values sampled from a Poisson distribution with a mean of 4 with a statistical variation of 2. If our distribution deviates from this it could indicate that the PTE value estimates are biased due to, for example, under- or overestimation of the uncertainty, or, for the cosmological tests, due to a disagreement with the GR-based predictions. We show a histogram of the PTE tests in bins of 10% in Fig. 7. For each bin in the histogram, the number of PTE values is within statistical expectations.
For an individual test, a PTE value of less than 0.05 is generally considered statistically significant as it has a change of only 5% to occur by chance (approximately the equivalent of 2 for the case of a Gaussian). However, when reporting multiple tests some outliers become more probable. For 40 tests having one outlier PTE value at or below 0.05 by chance has a probability of approximately 87% and having an outlier at or below 0.01 has a chance of approximately 33%. Among the reported PTE tests there are no outliers with PTE values below 1% or above 0.99%. The lowest PTE value across all PTE tests is 0.04 for a null test discussed in Section V.2.4. Therefore we conclude that the reported PTE values are fully consistent with statistical expectations indicating that our suite of null and consistency tests confirms the accuracy of our analysis pipeline.
V.3 Measurement results for cross-correlation of ACT and SDSS BOSS
@@
We show the measurements using ACT together with CMASS and LOWZ respectively. The top panels show the measurements with the baseline simulation-based error estimates. The white background range indicates the scales used for the cosmological analysis. The full range plotted is the extended validation range for which we also performed null and consistency tests. For comparison, the theory curve for our fiducial cosmology is also shown with dots indicating the binned measurement using the NaMaster coupling matrix. Note that this theory curve is not tuned to the amplitude of the measurement: the galaxy bias used is based on the amplitude of the angular auto-power spectrum of the galaxy data. In this work, we focus on constraining the statistic for which we directly use the measurements as part of the estimator, and the predicted theory values for are estimated from independent constraints on . In future work, the cross-correlation measurements we make available as part of this work can be consistently fitted with angular power spectra theory curves by varying the galaxy bias and all cosmological parameters. Specifically, we can visually see that for ACT CMASS the theory curve tuned to the measured galaxy auto-correlation amplitude appears slightly high compared to the cross-correlation measurement. This implies that when jointly fitting the cross-correlation one will likely find a lower best fit value, which quantifies the amplitude, than for the CMASS galaxy data alone. To evaluate whether this difference is significant or within statistical expectations careful theoretical modeling of the angular power spectra is needed in future work. We here provide the tested and validated cross-correlation measurements between ACT and CMASS that could be used for such an investigation. The ACT DR6 lensing data has already been analyzed in this way for combinations with galaxy data from unWISE [64] and similar investigations have been done for Planck lensing data [e.g. 72]. For this work, we focus on interpreting the measurement at linear scales through the statistic to interpret implications about gravity.
In the middle panels, the error for each bin is compared. The simulation-based estimate of the errors is closely consistent with the analytic estimate. This consistency also extends to the off-diagonal elements of the covariance as discussed in Section V.1.2. We also show the uncertainties based on a jackknife approach which have larger statistical uncertainty and are a systematic overestimate. The jackknife errors are consistent with the other estimates within statistical expectations.
@@
Based on the measured angular cross-power spectra we derive constraints on the statistic for the ACT DR6 CMB lensing data combined with CMASS and LOWZ in Section VI. In the following, we combine the measurement with results based on Planck CMB lensing data.
V.4 Combined cross-correlation measurement with ACT and Planck
@@
In Fig. 10 we show the results for combining the full baseline ACT SDSS and Planck SDSS angular cross-power spectra. We show the correlation with the CMASS sample on the left and with the LOWZ sample on the right. For each, we show the results using ACT CMB lensing discussed in Section V.3 and the results using Planck SDSS presented in Wenzl et al. [27]. We can visually see that the two measurements are closely consistent with each other, especially on large scales, which is expected since they are strongly correlated. Finally, we show the results for the minimum variance combination of the two measurements. In the bottom panels, we show the SNR for each measurement as well as the combination. We again additionally show the SNR estimates using the theory curves (dashed black lines) which represent a fair comparison between the measurements.
VI EG measurement
VI.1 Measurement approach for EG
@@
In Section VI.1.1 we discuss the creation of correctly correlated simulated maps and the estimate of the cross-covariance between and .
We use estimates of and for the full galaxy maps. To quantify the potential systematic caused by differences in galaxy bias between the full sample and the ACT DR6 overlap we constrain the galaxy bias difference and derive an estimate for our systematic error budget in Section VI.1.2.
When calculating the estimate using Eq. 14 we account for the full covariance of and , including their cross-covariance (see Section VI.1.1). For this, we calculate the ratio distribution888The probability distribution for the ratio of variables from two known distributions. of . Then in a second step, we calculate the ratio distribution of this estimate and resulting in a probability density function (PDF) for marginalized across the cosmological range or for individual bins . Notably, while we assume the measurements of and to be Gaussian distributed we do not assume the ratio to be Gaussian as the PDF is expected to be significantly asymmetric [27]. We report the median and the central confidence range of the PDF which generalize the best fit and uncertainty for the Gaussian case. For a full description of this calculation and a derivation of the PTE value tests to evaluate the results for consistency with scale independence () and for consistency with a predicted value () see Wenzl et al. [27].
VI.1.1 Cross-covariance estimation
@@
For our baseline result using ACT and SDSS BOSS, we leverage the set of 400 simulated map realizations to estimate the full covariance. As described in Section III.3, we have 400 simulated map realizations with the CMB lensing map and two galaxy maps for each with one galaxy map for the angular auto-power spectrum and the other for the angular cross-power spectrum where an additional redshift dependent weight was applied. For each realization, the three maps are correlated as expected for our fiducial cosmology and they contain the same noise level as our measurements. Using Eq. 28 we estimate the simulation-based covariances of and individually as well as the cross-covariance between them. For the full analytic covariance, we use NaMaster as described in Section III.3. For the covariance of we use a shot noise of and for we use as given by Eq. 24. For the cross-covariance, we use . We calculated the full simulation-based and analytic covariances. We find consistent results between the two methods within the statistical uncertainty of the simulation-based covariance. For the same bin, the correlation between the two measurements reaches around 33% for ACT + CMASS and 28% for ACT + LOWZ and drops off for smaller scales.
VI.1.2 Effective galaxy bias difference in ACT overlap
@@
We can perform a test for the ratio of the inferred between the overlap and the full sample. Assuming the underlying matter power spectra and redshift distribution are sufficiently similar for both the full sample and the overlap with ACT, this ratio directly tests if the effective linear galaxy bias is consistent between the two. By assuming that a constant linear bias describes the data well in our cosmological range, and therefore , we can put a constraint on the fractional difference in bias between the two . Since this approach avoids having to marginalize over uncertainty in the matter power spectrum this gives us a strong constraint on the fractional bias difference of the bias between the samples.
@@
The uniformity of the effective galaxy bias of the SDSS samples across the footprint permits us to combine our cross-correlation with the full auto-correlation and RSD analysis for the SDSS datasets [27] to measure . For completeness, we also present the result when the auto-correlation is limited to the region overlapping with ACT (see Section VI.3).
VI.2 Baseline EG results
@@
In Fig. 12 we show the measurement of as a function of scale. These are calculated using the estimator of Eq. 14, using the measured angular cross-power spectra between the ACT DR6 lensing map and the SDSS BOSS DR12 CMASS and LOWZ galaxy samples together with the angular auto-power spectra of the galaxy samples and from an RSD analysis of the galaxy samples. We account for the cross-covariance between the cross- and auto-power spectra and calculate the PDF for without assuming Gaussianity in the result. The Figure shows the median of the PDF for each bin in angular scale and the uncertainty is based on the central 68.27% confidence interval of the PDF. The constraints are shown for three different estimates of the uncertainty. The leftmost in orange are the results with the simulation-based covariance, in the middle in blue the constraints using the analytic covariance, and on the right in green the constraints using a jackknife-based covariance. In Section V.1.2 we described in detail how the simulation-based covariance and analytic covariance are consistent within statistical expectations. Here they give consistent results for within the statistical accuracy of the simulation-based covariance. The jackknife-based covariance serves as a consistency check. It is a systematic overestimate of the uncertainty of 20-30% and has large statistical uncertainty as it is based on only a small number of subsamples. Its advantage is that it is based on the data itself and could reveal additional sources of uncertainty not captured in the other methods. We here again find it to be consistent with the other methods within statistical expectations giving confidence to the accuracy of the approach. Finally Fig. 12 also shows the CDM prediction for reference. A key feature of the GR-based predictions for is that on linear scales the statistic is expected to be scale-independent. Therefore, before comparing the overall amplitude with the CDM-based prediction we can more generally test if the measurement is consistent with scale independence. We perform a PTE value test for consistency of the measurement with a scale-independent constant. The constant here is a free parameter not fixed to the CDM prediction. We find for ACT + CMASS and for ACT + LOWZ and therefore no statistical evidence for a deviation from scale-independence within the accuracy of the measurement. For ACT + CMASS 9 out of the 11 bins are below the GR prediction. However, the different bins are significantly correlated so to evaluate if this is statistically significant we combine the measurements under consideration of the full covariance.
@@
Since the measurement is consistent with a constant we combine the constraints across the cosmological analysis range into an overall constraint on . For this, we account for the full covariance between and (see Section VI.1.1). For the jackknife-based uncertainty estimate, there are not enough samples to invert this full covariance, therefore these combined constraints are calculated using the simulation-based and the analytic covariances. Fig. 13 shows the PDFs for the overall constraints. The statistic is measured for the effective redshift of the galaxy sample as given by Eq. 19, which results in for CMB lensing combined with CMASS and in for combinations with LOWZ. At these redshifts the expected values for the statistic for our fiducial cosmology assuming CDM are given by and . These predictions are shown as black vertical lines in Fig. 13. The Gaussian PDF for the predictions is too highly peaked to show in the same plot so the uncertainty for the predicted values is shown as a gray band instead. These predictions can be compared with the measurements based on ACT + CMASS and ACT + LOWZ. In each case, the PDF using the simulation-based covariance and the analytic covariance are shown. The results using either covariance are consistent. For the simulation-based covariance, the median is highlighted with a vertical orange line, and the central 68.27% confidence interval is shaded in orange. These result in and respectively. We note that the PDFs, especially for the LOWZ case, are significantly non-Gaussian and result in asymmetric uncertainties. The asymmetry is such that the error towards higher values is larger than towards lower values.
@@
We test the measured values for consistency with the CDM prediction resulting in for ACT + CMASS and for ACT + LOWZ. Both measurements are within statistical expectations, showing no statistical deviation from the prediction based on CDM. For ACT + CMASS the 68.27% confidence range does not include the prediction, i.e. the measurement differs by more than the equivalent of from the prediction. However, the deviation is not statistically significant with , meaning the 95% confidence interval, approximately equivalent to the error of a Gaussian, includes the prediction. Therefore within the statistical constraining power of the measurements, the data shows no evidence for a deviation from the CDM prediction. The results and PTE value tests are summarized in Table 2. In Section VI.3, we evaluate if these conclusions are sensitive to analysis choices. For an overview of values in alternative gravity scenarios see Pullen et al. [29].
@@
In addition to the statistical uncertainty we also estimate the systematic bias in the results. We combine in quadrature the conservative systematic bias estimates from galaxy bias evolution with redshift (Section II.2), the level at which we validated the analysis pipeline (Section V.2.1) and the upper limit on the effective galaxy bias difference between the full galaxy maps and the overlap with ACT DR6 lensing. We find systematic error budgets of 3% for ACT + CMASS and 4% for ACT + LOWZ. These are significantly smaller than the statistical uncertainties of the measurement (approximately 18% and 26% respectively) and therefore negligible when evaluating the consistency of the measurement with GR predictions.
VI.3 Sensitivity to analysis choices